# Qingdao 266237, China

### v1

@@ 0.F.1 Theoretical Analysis of Protocol $\Pi_{\mathsf{bssPMT}}$
  • Batch ssPMT The cost of this stage is the same as that in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (cf. Appendix 0.F.3).
    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
@@
    • Offline phase. The computation complexity of each party is $O(mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$ public-key operations. The communication complexity of each party is $O(\lambda mn)$. The round complexity is $O(1)$.
@@
    • Online phase. The computation complexity of each party is $O(mn)$ public-key operations. The communication complexity of each party is $O(\lambda mn)$. The round complexity is $O(m)$.
@@
    • Offline phase. The computation complexity of each party is $O(\gamma mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(\gamma n/t)$. The round complexity is $O(1)$.
    • Online phase. The computation complexity of each party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations. The communication complexity of each party is $O((\gamma+\lambda)mn)$. The round complexity is $O(\log\gamma+m)$.
  • 0.F.4 The Complete Analysis of LG Protocol
    The costs of each stage in LG protocol are calculated as follows.
    • mq-ssPMT. The cost of this stage mainly consists of three parts:
      • Offline phase. The computation complexity of each party is $O((T+l-\log n)mn\log((T+l-\log n)n))$. The communication complexity of each party is $O(t\lambda m\log(((T+l-\log n)n)/t))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity of each party is $O((T+l-\log n)mn)$. The communication complexity of each party is $O((T+l+\kappa-\log n)mn)$. The round complexity is $O(\log(l-\log n))$.
    • Random oblivious transfers. For $1\leq i<j\leq m$, $P_{i}$ and $P_{j}$ invoke ROT extension of size $B$. Overall, each party $P_{j}$ engages in $(m-1)$ independent instances of ROT extension of size $B$.
      • Offline phase. The computation complexity of each party is $O(mn\log n)$. The communication complexity of each party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
      • Online phase. The computation complexity of each party is $O(mn)$. The communication complexity of each party is $O(mn)$. The round complexity is $O(1)$.
      The costs of the remaining two steps, multi-party secret-shared shuffle and output reconstruction, are exactly the same as those in our SK-MPSU protocol.(cf. Appendix  0.F.2).
    • Total costs.
      • Offline phase. The computation complexity of each party is $O((T+l-\log n)mn\log((T+l-\log n)n)+m^{2}n\log mn)$. The communication complexity of each party is $O(t\lambda m\log(((T+l-\log n)n)/t)+\lambda m^{2}n(\log m+\log n))$. The round complexity is $O(1)$.
    Appendix 0.G Multi-Party Private-ID
    We generalize the two-party private-ID [BKM+20] to multi-party setting. Suppose there are $m$ parties, each possessing a set of $n$ elements. The multi-party private-ID functionality assigns a unique random identifier to each element across all input sets, ensuring that identical elements in different sets obtain the same identifier. Each party receives identifiers associated with its own input set, as well as identifiers associated with the union of all parties’ input sets. With multi-party private-ID, the parties can sort their private sets based on a global set of identifiers and perform desired private computations item by item, ensuring alignment of identical elements across their sets. The formal definition of the multi-party private-ID is depicted in Figure 18. We build a concrete multi-party private-ID protocol based on the DDH assumption (described in Figure 19) by extending the “distributed OPRF+PSU” paradigm [CZZ+24b] to multi-party setting.
@@ Appendix 0.G Multi-Party Private-ID
        For $1\leq i\leq m$, $P_{i}$ receives $\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}}$ for $j\in[n]$, then it computes $(\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}})^{-a_{i}k_{i}}=\mathsf{H}(x_{i}^{j})^{k_{1}\cdots k_{m}}$. We denote the set $\{\mathsf{H}(x_{i}^{1})^{k_{1}\cdots k_{m}},\mathsf{H}(x_{i}^{n})^{k_{1}\cdots k_{m}}\}$ as $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$, where each $r_{i}^{j}\in\mathbb{G}$.
@@
        The parties invoke $\Pi_{\mathsf{PK\text{-}MPSU}}$ where $P_{i}$ inputs $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$. $P_{1}$ receives the union $R^{*}=\bigcup_{i=1}^{m}Y_{i}$, and sends it to other parties.
@@
    • Correctness. The first two steps essentially realize a multi-party distributed OPRF protocol, where each party $P_{i}$ inputs a set $\{x_{i}^{1},\cdots,x_{i}^{n}\}$ and receives its own PRF key $k_{i}$ and the PRF values computed on its input set using all parties’ keys $k_{1},\cdots,k_{m}$, denoted as $\{\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{1}),\cdots,\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{n})\}$. In this case, $\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x)=\mathsf{H}(x)^{k_{1}\cdots k_{m}}$. Note that even if $m-1$ parties collude, there exist one exponent private to the adversary, ensuring that the PRF values remain pseudorandom (The proof is analogous to that of the DH-based OPRF, so we omit it here).
    • Security. Assuming $\mathsf{H}$ is a random oracle, the security of protocol follows immediately from the DDH assumption and the security of our PK-MPSU protocol.

### v4

@@ 0.E.4.1 Complete Analysis
  • Batch ssPMT The cost of this stage is the same as that in $\Pi_{\mathsf{SK\text{-}MPSU}}$ (cf. Appendix 0.E.4).
    • Offline phase. The computation complexity per party is $O(\gamma mn\log n)$. The communication complexity per party is $O(t\lambda m\log(\gamma n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(mn)$. The communication complexity per party is $O(\gamma mn)$. The round complexity is $O(\log\gamma)$.
@@
    • Offline phase. The computation complexity per party is $O(mn\log n)$. The communication complexity per party is $O(t\lambda m\log(n/t))$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(mn)$ public-key operations. The communication complexity per party is $O(\lambda mn)$. The round complexity is $O(1)$.
@@
    • Online phase. The computation complexity per party is $O(mn)$ public-key operations. The communication complexity per party is $O(\lambda mn)$. The round complexity is $O(m)$.
@@
    • Offline phase. The computation complexity per party is $O(\gamma mn\log n)$. The communication complexity per party is $O(t\lambda m\log(\gamma n/t)$. The round complexity is $O(1)$.
    • Online phase. The computation complexity per party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations. The communication complexity per party is $O((\gamma+\lambda)mn)$. The round complexity is $O(\log\gamma+m)$.
  • 0.E.4.2 Comparison with [GNT23]
    Appendix 0.F Multi-Party Private-ID
    It is natural to generalize the two-party private-ID [BKM+20] to multi-party setting. Suppose there are $m$ parties, each possessing a set of $n$ elements. The multi-party private-ID functionality assigns a unique random identifier to each element across all input sets, ensuring that identical elements in different sets obtain the same identifiers. Each party receives identifiers associated with its own input elements, as well as identifiers associated with the union of all parties’ input sets. With multi-party private-ID, the parties can sort their private sets based on a global set of identifiers and perform desired private computations item by item, ensuring alignment of identical elements across their sets. The formal definition of the multi-party private-ID is depicted in Figure 13. We build a concrete multi-party private-ID protocol based on the DDH assumption (described in Figure 14) by extending the “distributed OPRF+PSU” paradigm [CZZ+24b] to multi-party setting.
@@ Appendix 0.F Multi-Party Private-ID
        For $1\leq i\leq m$, $P_{i}$ receives $\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}}$, $j\in[n]$, and computes $(\mathsf{H}(x_{i}^{j})^{a_{i}k_{i+1}\cdots k_{m}k_{1}\cdots k_{i-1}})^{-a_{i}k_{i}}=\mathsf{H}(x_{i}^{j})^{k_{1}\cdots k_{m}}$. We denote $\{\mathsf{H}(x_{i}^{1})^{k_{1}\cdots k_{m}},\mathsf{H}(x_{i}^{n})^{k_{1}\cdots k_{m}}\}$ as $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$, where each $r_{i}^{j}\in\mathbb{G}$.
@@
        The parties invoke $\mathcal{F}_{\mathsf{mpsu}}$ where $P_{i}$ inputs $R_{i}=\{r_{i}^{1},\cdots,r_{i}^{n}\}$. $P_{1}$ receives the union $R^{*}=\bigcup_{i=1}^{m}Y_{i}$, and sends it to other parties.
@@
    Theorem 0.F.1
    Protocol $\Pi_{\mathsf{MPID}}$ securely implements $\mathcal{F}_{\mathsf{MPID}}$ against any semi-honest adversary corrupting $t<m$ parties in the $\mathcal{F}_{\mathsf{mpsu}}$-hybrid model, assuming the DDH assumption and $\mathsf{H}$ is a random oracle.
    • Correctness. The first two steps essentially realize a multi-party distributed OPRF protocol, where each party $P_{i}$ inputs a set $\{x_{i}^{1},\cdots,x_{i}^{n}\}$ and receives its own PRF key $k_{i}$ along with the PRF values computed on its input set using all parties’ keys $k_{1},\cdots,k_{m}$, denoted as $\{\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{1}),\cdots,\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x_{i}^{n})\}$. In Figure 14, each party $P_{i}$ first masks the hash value of each element by raising it to the power of $a_{i}$. Then these masked values are then cycled among the parties, with each party raising them to the power of its key. Finally, $P_{i}$ unmasks the values and raises them to the power of its won key $k_{i}$, resulting in the distributed PRF values $\mathsf{PRF}_{k_{1},\cdots,k_{m}}(x)=\mathsf{H}(x)^{k_{1}\cdots k_{m}}$. Even if $m-1$ parties collude, there still exists one exponent private to the adversary, ensuring that the PRF values remain pseudorandom according to the DDH assumption.
      In the step 3, the parties invoke an MPSU protocol on their distributed PRF values, so than they can securely receive all PRF values on the union of input sets.
    • Security. Assuming $\mathsf{H}$ is a random oracle, the security of protocol follows immediately from the DDH assumption and the security of our PK-MPSU.
      Modern DH-based cryptosystems often use elliptic curves for the underlying cyclic group, due to their compact size and better efficiency at a given level of security. In this context, the first two steps of the protocol have relatively low communication costs, meanwhile, the element space of the MPU protocol in step 3 is EC points, which aligns perfectly with the element space of our PK-MPSU. This alignment allows us to instantiate the MPSU functionality using our PK-MPSU, leveraging its lowest communication costs. As a result, the entire multi-party private-ID protocol achieves extremely low communication costs, making it well-suited for deployment in bandwidth-restricted environments.
      Another benefit of instantiating the MPSU functionality with our PK-MPSU is to achieve linear computation and communication complexity. Since the multi-party distributed OPRF phase has linear computation and communication, by integrating our PK-MPSU, the entire multi-party private-ID protocol can maintain the linear complexities throughout.